a)Three resistors, each having a resistance, R, are connected in parallel to a 1.50 V battery. If the resistors dissipate a total power of 3.00 W, what is the value of R?

b) Calculate the charge stored on a 2.0 nF capacitor connected to a 12.0 V battery.
Answer #1

Equivalent resistance of the parallel combination Rp is

1/Rp=1/R+1/R+1/R=3/R

Rp=R/3

Power P=V^{2}/Rp

3=1.5^{2}/(R/3)            (V=1.4volt, P=3watts)

R=3X1.5^{2}/3=2.25\Omega

b)charge q=CV

C=capacitance=2nF=2x10^{-9}F

V=voltage=12V

q=CV=2x10^{-9}x12=24nC=2.4X10^{-8}C=charge